1/2x+ (-1/2)y=1/2
(-1/3)x+(-3/4)y=2
I divided everything by 1/2 and you get x-y=1 for the first equation.
the second equation i times everything by 12 and get (-4)x + (-9)y = 24
multiply the first equation by -9, so we get -9x +9y = -9.
(-9)x + (9)y = -9
(-4)x + (-9)y = 24
I get (-15/13, -28/13)
But it is wrong, what did i do wrong? Thanks for the help!

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

1/2x+ (-1/2)y=1/2
(-1/3)x+(-3/4)y=2
-----------------------------
Multiply thru 1st by 2
Multiply thru 2nd by -12
-------------------------------
x - y = 1
4x +9y = -24
---
Multiply thru 1st by 4
----------------------------
4x - 4y = 4
4x + 9y = -24
----------------
Subtract 1st from 2nd and solve for "y":
13y = -28
y = -28/13
-----
Substitute into x - y = 1 to solve for "x":
x +(28/13) = 1
x = (13/13)-(28/13)
x = -15/13
==================
Cheers,
Stan H.

Answer by solver91311(24713)


You didn't do anything wrong. Find the person that told you that your answer is wrong and bet them a lot of money that you are actually right. Because you are. Given that you didn't make any mistakes writing the equations down in the first place.

John